OpenStudy (anonymous):

integrate(6x^5dx/x^6+1)dx

OpenStudy (anonymous):

Is this the question?\[\int\limits{6x^5 \over x^6+1}dx\]

OpenStudy (anonymous):

Yes!!! lol

OpenStudy (anonymous):

we need to do u substitution

OpenStudy (anonymous):

\[\int\limits\limits{6x^5 \over x^6+1}dx\]\[u=x^6+1\]\[du=6x^5 \space dx\]Hey look..! we have that just sitting in the numerator. So we can replace it with just du.\[\int\limits{du \over u}\]

OpenStudy (anonymous):

\[\int\limits{1 \over u}du=\ln|u|+C\]Don't forget to replace "u"\[\ln|x^6+1|+C\]

OpenStudy (anonymous):

can i just say your amazing

OpenStudy (anonymous):

Oh stop it

OpenStudy (anonymous):

lol i suck at math .. i have another problem i was wondering if you can help me please

OpenStudy (anonymous):

ok, use the equation button to post your question this time. type "int" for the integral sign then the rest

OpenStudy (anonymous):

ok dont go lol

OpenStudy (anonymous):

I only have 8 mins hurry!!

OpenStudy (anonymous):

wait ok lol

OpenStudy (anonymous):

int(2/x^2-3/x+e^3x)dx

OpenStudy (anonymous):

sorry idk how to use equation thing

OpenStudy (anonymous):

\[\int\limits{2 \over x^2}-{3 \over x}+e^{3x} \space dx\]?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

\[\int\limits\limits{2 \over x^2}-{3 \over x}+e^{3x} \space dx\]we can write this is separate integrals\[\int\limits\limits{2 \over x^2}dx-\int\limits{3 \over x}dx+\int\limits e^{3x} \space dx\]write powers in the negative fashion\[\int\limits\limits{2x^{-2}}dx-\int\limits{3 \over x}dx+\int\limits e^{3x} \space dx\]You can pull constants out in front\[2\int\limits\limits\limits{x^{-2}}dx-3\int\limits\limits{dx \over x}+\int\limits\limits e^{3x} \space dx\]\[-4x^{-1}-3\ln|x|+{1 \over 3}e^{3x}+C\]You have to u substitute the 3x in the last one. I g2g. Good luck

OpenStudy (anonymous):

-4x^-1 is supposed to be positive. My fault. 2x^-2, -2 comes out in front, add 1 to the power, all over the power. This will put it over -1 and reverse the sign. I hope you read this in time

OpenStudy (anonymous):

Thank you soooo much you are amazing